# Pauli Hamiltonians with an Aharonov-Bohm flux

### v1

@@ 3.3. Spectral and scattering properties
$\left(R_{\mathrm{P},\pm}^{(\mathrm{F})}(\lambda)\bm{\psi}\right)_{\!s}\!(r,\theta)=\!\int^{\infty}_{0}\!\!\!\!\mathrm{d}r^{\prime}\,r^{\prime}\!\int^{2\pi}_{0}\!\!\!\!\mathrm{d}\theta^{\prime}\,\sum_{\ell\in\mathbb{Z}}I_{|\ell+\alpha|}\big{(}\mp i\sqrt{\lambda}\,(r\wedge r^{\prime})\big{)}\,K_{|\ell+\alpha|}\big{(}\mp i\sqrt{\lambda}\,(r\land r^{\prime})\big{)}\,\psi_{s}(r^{\prime},\theta^{\prime})\,\tfrac{e^{i\ell(\theta-\theta^{\prime})}}{2\pi}\\ =\tfrac{i}{4}\int^{\infty}_{0}\!\!\!\!\mathrm{d}r^{\prime}\,r^{\prime}\!\int^{2\pi}_{0}\!\!\!\!\mathrm{d}\theta^{\prime}\,\sum_{\ell\in\mathbb{Z}}J_{|\ell+\alpha|}\big{(}\!\pm\!\sqrt{\lambda}\,(r\wedge r^{\prime})\big{)}\,H^{(1)}_{|\ell+\alpha|}\big{(}\!\pm\!\sqrt{\lambda}\,(r\wedge r^{\prime})\big{)}\,\psi_{s}(r^{\prime},\theta^{\prime})\,e^{i\ell(\theta-\theta^{\prime})}\,.$
@@
$f^{(\mathrm{F},\pm)}_{\mathbf{k}}(\theta)=\tfrac{e^{\pm i\frac{\pi}{4}}}{(2\pi)^{3/2}\sqrt{k}}\sum_{\ell\in\mathbb{Z}}\left(e^{\pm i\pi|\ell+\alpha|}-e^{\pm i\pi|\ell|}\right)e^{i\ell(\theta-\omega_{\pm})}\,.$ (3.14)
@@
$\int_{0}^{\infty}\!\!\!\mathrm{d}r\;r\;J_{|\ell+\alpha|}(k\,r)\,J_{|\ell+\alpha|}(k^{\prime}r)=\lim_{\zeta\to+\infty}\int_{0}^{\infty}\!\!\!\mathrm{d}r\;\frac{\zeta^{2}\,r}{\zeta^{2}+r^{2}}\;J_{|\ell+\alpha|}(k\,r)\,J_{|\ell+\alpha|}(k^{\prime}r)\\ =\lim_{\zeta\to+\infty}\zeta^{2}\,I_{|\ell+\alpha|}\big{(}\zeta\,(k\wedge k^{\prime})\big{)}\,K_{|\ell+\alpha|}\big{(}\zeta\,(k\vee k^{\prime})\big{)}=\tfrac{1}{2k}\,\delta(k^{\prime}\!-k)=\delta\big{(}(k^{\prime})^{2}-k^{2}\big{)}\,.$
@@
Let $\Theta\in\mathrm{M}_{4,\mathrm{Herm}}(\mathbb{C})$. Then, there exists a (normal-(((possibly empty)normal-))) discrete set $\mathrm{e}_{+}^{(\Theta)}$, with #$\#\big{[}\mathrm{e}_{+}^{(\Theta)}\big{]}\leqslant 4$, such that for any $\lambda\in\mathbb{R}^{+}\setminus\mathrm{e}_{+}^{(\Theta)}$ the limits
@@
$\big{(}\mathcal{G}_{\pm}(\lambda)\,\mathbf{q}\big{)}_{s}(r,\theta)=\lim_{\varepsilon\to 0^{+}}\big{(}\mathcal{G}(\lambda\pm i\varepsilon)\,\mathbf{q}\big{)}_{s}(r,\theta)=\tfrac{i\pi}{2}\,{\textstyle\sum_{\ell}}\;q_{s}^{(\ell)}\;\big{(}\!\mp\!\sqrt{\lambda}\,\big{)}^{|\ell+\alpha|}\,H^{(1)}_{|\ell+\alpha|}\big{(}\!\mp\!\sqrt{\lambda}\,r\big{)}\,\tfrac{e^{i\ell\theta}}{\sqrt{2\pi}}\,,$ (3.21)
@@
$\big{(}\bm{\varphi}^{(\Theta,\pm)}_{(s,\mathbf{k})}\big{)}_{\!s^{\prime}}(r,\theta)=\tfrac{1}{2\pi}\,\sum_{\ell\in\mathbb{Z}}e^{i\ell(\theta-\omega_{\pm})\pm i\frac{\pi}{2}\,|\ell+\alpha|}\,J_{|\ell+\alpha|}(k\,r)\,\delta_{s\,s^{\prime}}\,\\ +\tfrac{i}{4}\sum_{\ell,\ell^{\prime}\in\{0,-1\}}\!\Big{\{}\big{[}\Lambda_{\pm}(k^{2})+\Theta\big{]}^{-1}\Big{\}}_{s^{\prime}\!,s}^{\ell^{\prime}\!,\ell}\,e^{i(\ell^{\prime}\theta-\ell\omega_{\pm})}\;(\pm ik)^{|\ell+\alpha|}\,(\mp k)^{|\ell^{\prime}+\alpha|}\,H^{(1)}_{|\ell^{\prime}+\alpha|}(\mp k\,r)\,.$
@@
$f^{(\Theta,\pm)}_{(s,s^{\prime}),\mathbf{k}}=\tfrac{e^{\pm i\frac{\pi}{4}}}{(2\pi)^{3/2}\sqrt{k}}\sum_{\ell\in\mathbb{Z}}\left(e^{\pm i\pi|\ell+\alpha|}-e^{\pm i\pi|\ell|}\right)e^{i\ell(\theta-\omega_{\pm})}\delta_{s\,s^{\prime}}\,\\ +\tfrac{i\pi\,e^{\pm i\frac{\pi}{4}}}{(2\pi)^{3/2}\sqrt{k}}\sum_{\ell,\ell^{\prime}\in\{0,-1\}}\!\Big{\{}\big{[}\Lambda_{\pm}(k^{2})+\Theta\big{]}^{-1}\Big{\}}_{s^{\prime}\!,s}^{\ell^{\prime}\!,\ell}\,e^{i(\ell^{\prime}\theta-\ell\omega_{\pm})}\;(\pm ik)^{|\ell+\alpha|+|\ell^{\prime}+\alpha|}\,.$ (3.23)
@@
$\psi_{s}(r,\theta)=\sum_{\ell\in\mathbb{Z}}\frac{e^{i\ell\theta}}{\sqrt{2\pi}}\left[c_{s}^{(\ell)}\,r^{-|\ell+\alpha|}+d_{s}^{(\ell)}\,r^{|\ell+\alpha|}\right],$ (3.24)
@@
$c_{s}^{(\ell)}=\left\{\!\begin{array}[]{ll}\frac{\Gamma(|\ell+\alpha|)}{2^{1-|\ell+\alpha|}}\,q_{s}^{(\ell)}&\mbox{for $\ell\in\{0,-1\}$}\,,\\ 0&\mbox{for $\ell\in\mathbb{Z}\setminus\{0,-1\}$}\,,\end{array}\right.\qquad\mbox{for some $\mathbf{q}\in\ker\big{[}\Lambda(0)+\Theta\big{]}$}\,,$
@@ 3.5. Comparison with the Dirac operator
Consider any of the self-adjoint extensions $H_{\mathrm{P}}^{(\beta)}$, and take $\bm{\psi}\in\mathscr{D}(H_{\mathrm{P}}^{(\beta)})$ as in (2.13). We first prove that if we additionally require that $\bm{\psi}\in\mathscr{D}\big{(}H_{\mathrm{D}}^{\gamma})^{2}\big{)}$, for some $\gamma\in[0,2\pi)$, then we must have $\bm{\psi}\in\mathscr{D}(H_{\mathrm{P}}^{(\mathrm{F})})$. In other words, our aim is to prove that
$\mathscr{D}\big{(}H_{\mathrm{P}}^{(\beta)}\big{)}\cap\mathscr{D}\big{(}(H_{\mathrm{D}}^{\gamma})^{2}\big{)}\subseteq\mathscr{D}\big{(}H_{\mathrm{P}}^{(\mathrm{F})}\big{)}\,.$
@@
$\mathscr{D}\big{(}H_{\mathrm{P}}^{(\mathrm{F})}\big{)}\subseteq\mathscr{D}\big{(}(H_{\mathrm{D}}^{\gamma})^{2}\big{)}\,,\quad\mbox{for any $\gamma\in[0,2\pi)$.}$ (3.27)
@@
$\bm{\psi}\in\mathscr{D}\big{(}H_{\mathrm{D}}^{\gamma}\big{)}\,,\qquad H_{\mathrm{D}}^{\gamma}\bm{\psi}\in\mathscr{D}\big{(}H_{\mathrm{D}}^{\gamma}\big{)}\,.$ (3.28)
We fix the first condition exploiting the characterization of the domain through boundary conditions at the origin, as in (2.67). Recalling (2.8) and (2.9), we observe that, since $\bm{\phi}_{\lambda}\in\mathscr{D}\big{[}Q^{(\mathrm{F})}_{P}\big{]}$, the spinor vanishes at the origin, so that, for $s\in\{\uparrow,\downarrow\}$,
@@
We now need to impose the second condition in (3.28), namely $H_{\mathrm{D}}^{\gamma}\bm{\psi}\in\mathscr{D}(H_{\mathrm{D}}^{\gamma})$. We preliminarily observe that $\phi_{\lambda}$ can be rewritten as
@@
$H_{\mathrm{D}}=\begin{pmatrix}0&e^{-i\theta}(-i\partial_{r}-\frac{\partial_{\theta}+i\alpha}{r})\\ e^{i\theta}(-i\partial_{r}+\frac{\partial_{\theta}+i\alpha}{r})&0\end{pmatrix}\,.$
@@
$(H_{\mathrm{D}}\bm{\phi}_{\lambda})_{\uparrow}=\sum_{k\in\mathbb{Z}}\left[\left(-i\partial_{r}-i\tfrac{(k+\alpha)}{r}\right)K_{|k+\alpha|}(\lambda r)\right]\int^{r}_{0}\,\mathrm{d}r^{\prime}r^{\prime}I_{|k+\alpha|}(\lambda r^{\prime})f_{\downarrow}^{k}(r^{\prime})\>\tfrac{e^{i(k-1)\theta}}{\sqrt{2\pi}}\\ +\sum_{k\in\mathbb{Z}}\left[\left(-i\partial_{r}-i\tfrac{(k+\alpha)}{r}\right)I_{|k+\alpha|}(\lambda r)\right]\int^{+\infty}_{r}\,\mathrm{d}r^{\prime}r^{\prime}K_{|k+\alpha|}(\lambda r^{\prime})f_{\downarrow}^{k}(r^{\prime})\>\tfrac{e^{i(k-1)\theta}}{\sqrt{2\pi}}\,,$ (3.33)
@@
$(H_{\mathrm{D}}\bm{\phi}_{\lambda})_{\downarrow}=\sum_{k\in\mathbb{Z}}\left[\left(-i\partial_{r}+i\tfrac{(k+\alpha)}{r}\right)K_{|k+\alpha|}(\lambda r)\right]\int^{r}_{0}\,\mathrm{d}r^{\prime}r^{\prime}I_{|k+\alpha|}(\lambda r^{\prime})f_{\uparrow}^{k}(r^{\prime})\>\tfrac{e^{i(k+1)\theta}}{\sqrt{2\pi}}\\ +\sum_{k\in\mathbb{Z}}\left[\left(-i\partial_{r}+i\tfrac{(k+\alpha)}{r}\right)I_{|k+\alpha|}(\lambda r)\right]\int^{+\infty}_{r}\,\mathrm{d}r^{\prime}r^{\prime}K_{|k+\alpha|}(\lambda r^{\prime})f_{\uparrow}^{k}(r^{\prime})\>\tfrac{e^{i(k+1)\theta}}{\sqrt{2\pi}}\,.$ (3.34)
@@
with the $\bm{\xi}_{\pm}$ as in (2.60). By the previous remarks, checking that $H_{\mathrm{D}}^{\gamma}\bm{\psi}\in\mathscr{D}(H_{\mathrm{D}}^{\gamma})$ leads to
$c^{\uparrow}_{-\alpha}\big{(}H_{\mathrm{D}}^{\gamma}\bm{\psi}\big{)}=0\,,\qquad c^{\downarrow}_{\alpha-1}\big{(}H_{\mathrm{D}}^{\gamma}\bm{\psi}\big{)}=0,$
@@
$\left.\big{(}H_{\mathrm{D}}^{\gamma}\big{)}^{2}\right|_{\mathscr{D}(H_{\mathrm{P}}^{(\mathrm{F})})}=H_{\mathrm{P}}^{(\mathrm{F})},$
and thus since $H_{\mathrm{P}}^{(\mathrm{F})}$ and $(H_{\mathrm{D}}^{\gamma})^{2}$ are both self-adjoint, they must coincide. To this aim, it is useful to recall the following properties of Bessel functions [GR07, Eq. 8.486]
@@
$\big{(}H_{\mathrm{D}}^{\gamma}\big{)}^{2}\bm{\phi}=H_{\mathrm{D}}^{2}\bm{\phi}=H_{\mathrm{P}}\bm{\phi}=H_{\mathrm{P}}^{(\mathrm{F})}\bm{\phi}\,,$

### v3

@@ 3.3. Spectral and scattering properties
$\left(R_{\mathrm{P},\pm}^{(\mathrm{F})}(\lambda)\bm{\psi}\right)_{\!s}\!(r,\vartheta)=\!\int^{\infty}_{0}\!\!\!\!\mathrm{d}r^{\prime}\,r^{\prime}\!\int^{2\pi}_{0}\!\!\!\!\mathrm{d}\vartheta^{\prime}\,\sum_{\ell\in\mathbb{Z}}I_{|\ell+\alpha|}\big{(}\mp i\sqrt{\lambda}\,(r\land r^{\prime})\big{)}\,K_{|\ell+\alpha|}\big{(}\mp i\sqrt{\lambda}\,(r\lor r^{\prime})\big{)}\,\psi_{s}(r^{\prime},\vartheta^{\prime})\,\tfrac{e^{i\ell(\vartheta-\vartheta^{\prime})}}{2\pi}\\ =\tfrac{i}{4}\int^{\infty}_{0}\!\!\!\!\mathrm{d}r^{\prime}\,r^{\prime}\!\int^{2\pi}_{0}\!\!\!\!\mathrm{d}\vartheta^{\prime}\,\sum_{\ell\in\mathbb{Z}}J_{|\ell+\alpha|}\big{(}\!\pm\!\sqrt{\lambda}\,(r\land r^{\prime})\big{)}\,H^{(1)}_{|\ell+\alpha|}\big{(}\!\pm\!\sqrt{\lambda}\,(r\lor r^{\prime})\big{)}\,\psi_{s}(r^{\prime},\vartheta^{\prime})\,e^{i\ell(\vartheta-\vartheta^{\prime})}\,.$
@@
$f^{(\mathrm{F},\pm)}_{\mathbf{k}}(\vartheta)=\tfrac{e^{\pm i\frac{\pi}{4}}}{(2\pi)^{3/2}\sqrt{k}}\sum_{\ell\in\mathbb{Z}}\left(e^{\pm i\pi|\ell+\alpha|}-e^{\pm i\pi|\ell|}\right)e^{i\ell(\vartheta-\omega_{\pm})}\,.$ (3.14)
@@
$\int_{0}^{\infty}\!\!\!\mathrm{d}r\;r\;J_{|\ell+\alpha|}(k\,r)\,J_{|\ell+\alpha|}(k^{\prime}r)=\lim_{\zeta\to+\infty}\int_{0}^{\infty}\!\!\!\mathrm{d}r\;\frac{\zeta^{2}\,r}{\zeta^{2}+r^{2}}\;J_{|\ell+\alpha|}(k\,r)\,J_{|\ell+\alpha|}(k^{\prime}r)\\ =\lim_{\zeta\to+\infty}\zeta^{2}\,I_{|\ell+\alpha|}\big{(}\zeta\,(k\land k^{\prime})\big{)}\,K_{|\ell+\alpha|}\big{(}\zeta\,(k\lor k^{\prime})\big{)}=\tfrac{1}{2k}\,\delta(k^{\prime}\!-k)=\delta\big{(}(k^{\prime})^{2}-k^{2}\big{)}\,.$
@@
Let $\Theta\in\mathrm{M}_{4,\mathrm{Herm}}(\mathbb{C})$. Then, there exists a ((((possibly empty)))) discrete set $\mathrm{e}_{+}^{(\Theta)}$, with #$\#\big{[}\mathrm{e}_{+}^{(\Theta)}\big{]}\leqslant 4$, such that for any $\lambda\in\mathbb{R}^{+}\setminus\mathrm{e}_{+}^{(\Theta)}$ the limits
@@
$\big{(}\mathcal{G}_{\pm}(\lambda)\,\mathbf{q}\big{)}_{s}(r,\vartheta)=\lim_{\varepsilon\to 0^{+}}\big{(}\mathcal{G}(\lambda\pm i\varepsilon)\,\mathbf{q}\big{)}_{s}(r,\vartheta)=\tfrac{i\pi}{2}\,{\textstyle\sum_{\ell}}\;q_{s}^{(\ell)}\;\big{(}\!\mp\!\sqrt{\lambda}\,\big{)}^{|\ell+\alpha|}\,H^{(1)}_{|\ell+\alpha|}\big{(}\!\mp\!\sqrt{\lambda}\,r\big{)}\,\tfrac{e^{i\ell\vartheta}}{\sqrt{2\pi}}\,.$ (3.21)
@@
$\big{(}\bm{\varphi}^{(\Theta,\pm)}_{(s,\mathbf{k})}\big{)}_{\!s^{\prime}}(r,\vartheta)=\tfrac{1}{2\pi}\,\sum_{\ell\in\mathbb{Z}}e^{i\ell(\vartheta-\omega_{\pm})\pm i\frac{\pi}{2}\,|\ell+\alpha|}\,J_{|\ell+\alpha|}(k\,r)\,\delta_{s\,s^{\prime}}\,\\ +\tfrac{i}{4}\sum_{\ell,\ell^{\prime}\in\{0,-1\}}\!\Big{\{}\big{[}\Lambda_{\pm}(k^{2})+\Theta\big{]}^{-1}\Big{\}}_{s^{\prime}\!,s}^{\ell^{\prime}\!,\ell}\,e^{i(\ell^{\prime}\vartheta-\ell\omega_{\pm})}\;(\pm ik)^{|\ell+\alpha|}\,(\mp k)^{|\ell^{\prime}+\alpha|}\,H^{(1)}_{|\ell^{\prime}+\alpha|}(\mp k\,r)\,.$
@@
$f^{(\Theta,\pm)}_{(s,s^{\prime}),\mathbf{k}}=\tfrac{e^{\pm i\frac{\pi}{4}}}{(2\pi)^{3/2}\sqrt{k}}\sum_{\ell\in\mathbb{Z}}\left(e^{\pm i\pi|\ell+\alpha|}-e^{\pm i\pi|\ell|}\right)e^{i\ell(\vartheta-\omega_{\pm})}\delta_{s\,s^{\prime}}\,\\ +\tfrac{i\pi\,e^{\pm i\frac{\pi}{4}}}{(2\pi)^{3/2}\sqrt{k}}\sum_{\ell,\ell^{\prime}\in\{0,-1\}}\!\Big{\{}\big{[}\Lambda_{\pm}(k^{2})+\Theta\big{]}^{-1}\Big{\}}_{s^{\prime}\!,s}^{\ell^{\prime}\!,\ell}\,e^{i(\ell^{\prime}\vartheta-\ell\omega_{\pm})}\;(\pm ik)^{|\ell+\alpha|+|\ell^{\prime}+\alpha|}\,.$ (3.23)
@@
$\psi_{s}(r,\vartheta)=\sum_{\ell\in\mathbb{Z}}\frac{e^{i\ell\vartheta}}{\sqrt{2\pi}}\left[c_{s}^{(\ell)}\,r^{-|\ell+\alpha|}+d_{s}^{(\ell)}\,r^{|\ell+\alpha|}\right],$ (3.24)
@@
$c_{s}^{(\ell)}=\left\{\!\begin{array}[]{ll}\frac{\Gamma(|\ell+\alpha|)}{2^{1-|\ell+\alpha|}}\,q_{s}^{(\ell)}&\mbox{for $\ell\in\{0,-1\}$}\,,\vspace{0.15cm}\\ 0&\mbox{for $\ell\in\mathbb{Z}\setminus\{0,-1\}$}\,,\end{array}\right.\qquad\mbox{for some $\mathbf{q}\in\ker\big{[}\Lambda(0)+\Theta\big{]}$}\,,$
@@ 3.5. Comparison with the Dirac operator
Consider any of the self-adjoint extensions $H_{\mathrm{P}}^{(\beta)}$, and take $\bm{\psi}\in\mathscr{D}(H_{\mathrm{P}}^{(\beta)})$ as in (2.13). We first prove that if we additionally require that $\bm{\psi}\in\mathscr{D}\big{(}H_{\mathrm{D}}^{(\gamma)})^{2}\big{)}$, for some $\gamma\in[0,2\pi)$, then we must have $\bm{\psi}\in\mathscr{D}(H_{\mathrm{P}}^{(\mathrm{F})})$. In other words, our aim is to prove that
$\mathscr{D}\big{(}H_{\mathrm{P}}^{(\beta)}\big{)}\cap\mathscr{D}\big{(}(H_{\mathrm{D}}^{(\gamma)})^{2}\big{)}\subseteq\mathscr{D}\big{(}H_{\mathrm{P}}^{(\mathrm{F})}\big{)}\,.$
@@
$\mathscr{D}\big{(}H_{\mathrm{P}}^{(\mathrm{F})}\big{)}\subseteq\mathscr{D}\big{(}(H_{\mathrm{D}}^{(\gamma)})^{2}\big{)}\,,\quad\mbox{for any $\gamma\in[0,2\pi)$.}$ (3.27)
@@
$\bm{\psi}\in\mathscr{D}\big{(}H_{\mathrm{D}}^{(\gamma)}\big{)}\,,\qquad H_{\mathrm{D}}^{(\gamma)}\bm{\psi}\in\mathscr{D}\big{(}H_{\mathrm{D}}^{(\gamma)}\big{)}\,.$ (3.28)
We fix the first condition exploiting the characterization of the domain through boundary conditions at the origin, as in (2.69). Recalling (2.8) and (2.9), we observe that, since $\bm{\phi}_{\lambda}\in\mathscr{D}\big{[}Q^{(\mathrm{F})}_{\mathrm{P}}\big{]}$, the spinor vanishes at the origin, so that, for $s\in\{\uparrow,\downarrow\}$,
@@
We now need to impose the second condition in (3.28), namely $H_{\mathrm{D}}^{(\gamma)}\bm{\psi}\in\mathscr{D}(H_{\mathrm{D}}^{(\gamma)})$. We preliminarily observe that $\phi_{\lambda}$ can be rewritten as
@@
$H_{\mathrm{D}}=\begin{pmatrix}0&e^{-i\vartheta}(-i\partial_{r}-\frac{\partial_{\vartheta}+i\alpha}{r})\\ e^{i\vartheta}(-i\partial_{r}+\frac{\partial_{\vartheta}+i\alpha}{r})&0\end{pmatrix}\,.$
@@
$(H_{\mathrm{D}}\bm{\phi}_{\lambda})_{\uparrow}=\sum_{k\in\mathbb{Z}}\left[\left(-i\partial_{r}-i\tfrac{(k+\alpha)}{r}\right)K_{|k+\alpha|}(\lambda r)\right]\int^{r}_{0}\,\mathrm{d}r^{\prime}r^{\prime}I_{|k+\alpha|}(\lambda r^{\prime})f_{\downarrow}^{k}(r^{\prime})\>\tfrac{e^{i(k-1)\vartheta}}{\sqrt{2\pi}}\\ +\sum_{k\in\mathbb{Z}}\left[\left(-i\partial_{r}-i\tfrac{(k+\alpha)}{r}\right)I_{|k+\alpha|}(\lambda r)\right]\int^{+\infty}_{r}\,\mathrm{d}r^{\prime}r^{\prime}K_{|k+\alpha|}(\lambda r^{\prime})f_{\downarrow}^{k}(r^{\prime})\>\tfrac{e^{i(k-1)\vartheta}}{\sqrt{2\pi}}\,,$ (3.33)
@@
$(H_{\mathrm{D}}\bm{\phi}_{\lambda})_{\downarrow}=\sum_{k\in\mathbb{Z}}\left[\left(-i\partial_{r}+i\tfrac{(k+\alpha)}{r}\right)K_{|k+\alpha|}(\lambda r)\right]\int^{r}_{0}\,\mathrm{d}r^{\prime}r^{\prime}I_{|k+\alpha|}(\lambda r^{\prime})f_{\uparrow}^{k}(r^{\prime})\>\tfrac{e^{i(k+1)\vartheta}}{\sqrt{2\pi}}\\ +\sum_{k\in\mathbb{Z}}\left[\left(-i\partial_{r}+i\tfrac{(k+\alpha)}{r}\right)I_{|k+\alpha|}(\lambda r)\right]\int^{+\infty}_{r}\,\mathrm{d}r^{\prime}r^{\prime}K_{|k+\alpha|}(\lambda r^{\prime})f_{\uparrow}^{k}(r^{\prime})\>\tfrac{e^{i(k+1)\vartheta}}{\sqrt{2\pi}}\,.$ (3.34)
@@
with the $\bm{\xi}_{\pm}$ as in (2.62). By the previous remarks, checking that $H_{\mathrm{D}}^{(\gamma)}\bm{\psi}\in\mathscr{D}(H_{\mathrm{D}}^{(\gamma)})$ leads to
$c^{\uparrow}_{-\alpha}\big{(}H_{\mathrm{D}}^{(\gamma)}\bm{\psi}\big{)}=0\,,\qquad c^{\downarrow}_{\alpha-1}\big{(}H_{\mathrm{D}}^{(\gamma)}\bm{\psi}\big{)}=0,$
@@
$\left.\big{(}H_{\mathrm{D}}^{(\gamma)}\big{)}^{2}\right|_{\mathscr{D}(H_{\mathrm{P}}^{(\mathrm{F})})}=H_{\mathrm{P}}^{(\mathrm{F})},$
and thus since $H_{\mathrm{P}}^{(\mathrm{F})}$ and $(H_{\mathrm{D}}^{(\gamma)})^{2}$ are both self-adjoint, they must coincide. To this aim, it is useful to recall the following properties of Bessel functions [GR07, Eq. 8.486]
@@
$\big{(}H_{\mathrm{D}}^{(\gamma)}\big{)}^{2}\bm{\phi}=H_{\mathrm{D}}^{2}\bm{\phi}=H_{\mathrm{P}}\bm{\phi}=H_{\mathrm{P}}^{(\mathrm{F})}\bm{\phi}\,,$
@@
Proof of Proposition 2.27.
Proceeding as in the proof of Proposition 2.21, it can be checked that the only distributional solutions of the zero-energy equation $H_{\mathrm{D}}\bm{\psi}=0$ are of the form
$\bm{\psi}(r,\vartheta)=\begin{pmatrix}\sum_{\ell\in\mathbb{Z}}d_{\uparrow}^{(\ell)}\,\frac{e^{i\ell\vartheta}}{\sqrt{2\pi}}\,r^{\ell+\alpha}\vspace{0.1cm}\\ \sum_{\ell\in\mathbb{Z}}d_{\downarrow}^{(\ell)}\,\frac{e^{i\ell\vartheta}}{\sqrt{2\pi}}\,r^{-(\ell+\alpha)}\end{pmatrix}\,,$ (3.38)
with suitable coefficients $\mathbf{d}\in\mathbb{C}^{4}$. The condition of uniform boundedness at infinity forces us to fix $d_{\uparrow}^{(\ell)}=0$ for $\ell\geqslant 0$ and $d_{\downarrow}^{(\ell)}=0$ for $\ell\leqslant-1$. On the other hand, the condition $\bm{\psi}\in L^{2}_{\mathrm{loc}}(\mathbb{R}^{2})$ entails $d_{\uparrow}^{(\ell)}=0$ for $\ell\leqslant-2$ and $d_{\downarrow}^{(\ell)}=0$ for $\ell\geqslant 1$. Summing up, the only admissible zero-energy resonances have the form
$\bm{\psi}(r,\vartheta)=\begin{pmatrix}d_{\uparrow}^{(-1)}\,\frac{e^{-i\vartheta}}{\sqrt{2\pi}}\,\frac{1}{r^{1-\alpha}}\vspace{0.1cm}\\ d_{\downarrow}^{(0)}\,\frac{1}{\sqrt{2\pi}}\,\frac{1}{r^{\alpha}}\end{pmatrix}\,.$ (3.39)
Then, it can be easily checked that the boundary conditions associated to $\mathscr{D}(H_{\mathrm{D}}^{(\gamma)})$, see (2.69), are verified if and only if
$d_{\uparrow}^{(-1)}=i\cot(\gamma/2)\,\tfrac{2^{1-2\alpha}\Gamma(1-\alpha)}{\Gamma(\alpha)}\,d_{\downarrow}^{(0)}\,,$
which concludes the proof. ∎